OpenStudy (animalover02):

I WILL MEDAL!!!!!!!!!!!!! Anita and joelle bowl together and there combined total score for one game was 425 points. Anita's score was 40 less than twice joelle's. what were there scores? A. joelle: 155 , Anita: 270 B. joelle: 180 , Anita: 245 C. joelle: 270 , Anita: 155 D. joelle: 140 , Anita: 240

OpenStudy (wanell):

x + y = 425 --> x = 425 - y now sub 425 - y in for x y = 2x - 40 y = 2(425 - y) - 40 -- distribute through the parenthesis y = 850 - 2y - 40 2y + y = 850 - 40 3y = 810 y = 810/3 y = 270 now sub 270 in for y in either of the equations to find x x + y = 425 x + 270 = 425 x = 425 - 270 x = 155 anita's score (y) = 270 joelle's score (x) = 155

OpenStudy (wanell):

does this help a little

OpenStudy (animalover02):

its c

OpenStudy (animalover02):

correct?

OpenStudy (acannell):

x + y = 425 --> x = 425 - y now sub 425 - y in for x y = 2x - 40 y = 2(425 - y) - 40 -- distribute through the parenthesis y = 850 - 2y - 40 2y + y = 850 - 40 3y = 810 y = 810/3 y = 270 now sub 270 in for y in either of the equations to find x x + y = 425 x + 270 = 425 x = 425 - 270 x = 155 anita's score (y) = 270 joelle's score (x) = 155

OpenStudy (wanell):

your welcome

OpenStudy (acannell):

Its A

OpenStudy (wanell):

we have the same work she just have to look at it

OpenStudy (animalover02):

what?

OpenStudy (acannell):

Oh lol

OpenStudy (acannell):

Its A

OpenStudy (animalover02):

ok